# MAINS: A Magnetic Field Aided Inertial Navigation System for Indoor Positioning

### 05 Dec 2023 v1

@@ II-B Magnetic Field Modeling
The magnetic field is a three-dimensional vector field whose properties are described by Maxwell’s equations. Let $M(r;\theta)$ be a model of the magnetic field at the location $r\in\mathbb{R}^{3}$, parameterized by the parameter $\theta$. When there is no free current in the space $\Omega$, the magnetic-field model $M(r;\theta)$ should fulfill
$\nabla_{r}\times M(r;\theta)$ $=0,$ (3a)
$\nabla_{r}\cdot M(r;\theta)$ $=0,$ (3b)
for all $r\in\Omega$. A polynomial magnetic field model, which fulfills (II-B), is given by [23]
$M(r;\theta)=\Phi(r)\theta.$ (4)
@@
Here $\Phi(r)\in\mathbb{R}^{3\times\kappa}$ is the regression matrix defined in [23] and $\theta\in\mathbb{R}^{\kappa}$ is the coefficient of the polynomial model; for a $n^{\text{th}}$ order polynomial the model has $\kappa$ = dim($\theta$) = $n^{2}+4n+3$ unknown parameters[2]. Note that the model (4) can be defined in either the body frame or navigation frame. Within this paper, it will be defined in the body frame. Next, a procedure for transforming the model from the $b_{k}$ frame to the $b_{k+1}$ frame will be presented.
@@
Let the magnetic field model (4) be associated with the $b$-frame, which means $M(r;\theta)$ accepts locations $r$ expressed in $b$-frame and outputs magnetic field vector in the same frame. Then $M_{k}(r;\theta_{k})$ denotes the magnetic field model associated with the $b_{k}$-frame, i.e.,
$M_{k}(r^{b_{k}};\theta_{k})=\Phi(r^{b_{k}})\theta_{k}.$ (5)
When the body frame moves, the coefficient $\theta_{k}$ evolves over time along with the changes in the body frame. Let the relative body frame change from $b_{k}$ to $b_{k+1}$ be encoded by
$\psi_{k}=\begin{bmatrix}\Delta p_{k}\\ \Delta\phi_{k}\end{bmatrix},$ (6)
where ${\Delta p}_{k}\in\mathbb{R}^{3}$ and $\Delta\phi_{k}\in[0,2\pi]^{3}$ denote the translation and orientation change from the body frame $b_{k}$ to $b_{k+1}$, respectively. Then the dynamics of $\theta_{k}$ can be found by expressing the magnetic field vector at a given location with the models at two consecutive times and aligning them in the same frame. The basic idea is illustrated in Fig. 3, from which the following two equalities can be identified
$M_{k+1}(r^{b_{k+1}};\theta_{k+1})$ $=R_{b_{k}}^{b_{k+1}}M_{k}(r^{b_{k}};\theta_{k}),$ (7a)
$r^{b_{k}}$ $=(R_{b_{k}}^{b_{k+1}})^{\top}r^{b_{k+1}}+{\Delta p}_{k}.$ (7b)
Here $R_{b_{k}}^{b_{k+1}}$ denotes the rotation matrix that rotates a vector from the $b_{k}$-frame to the $b_{k+1}$-frame. The rotation matrix $R_{b_{k}}^{b_{k+1}}$ and translation $\Delta p_{k}$ are given by
$R_{b_{k}}^{b_{k+1}}=\left(\text{exp}(\left[\Delta\phi_{k}\right]_{\times})\right)^{\top}=\left(\text{exp}(\left[\omega_{k}T_{s}\right]_{\times})\right)^{\top},$ (8a)
$\Delta p_{k}=R_{b_{k}}^{n^{\top}}\left(v_{k}^{n}T_{s}+(R^{n}_{b_{k}}s_{k}+g^{n})\frac{T_{s}^{2}}{2}\right).$ (8b)
@@
Here $[\cdot]_{\times}$ is an operator that maps a vector in $\mathbb{R}^{3}$ to a skew-symmetric matrix such that $\left[a\right]_{\times}b=a\times b$.
$x$$y$$b_{k}$$\Omega$$x^{\prime}$$y^{\prime}$$b_{k+1}$$\Delta\phi_{k}$ $\Delta p_{k}$ $r^{b_{k}}$ $r^{b_{k+1}}$
@@
Figure 3: A 2D illustration of the geometric relationship between the body frames at two consecutive times. The applicable region $\Omega$ of the magnetic field model at time $k$ is in blue, and the black dot indicates the location where the two models output the corresponding magnetic field in their own coordinate frames.
@@
Substituting the generic magnetic model in (II-B) with the proposed polynomial model (5) yields the equality
@@
Note that for a given $r^{b_{k+1}}$, $\theta_{k}$ and $\psi_{k}$, (9) represents 3 linear equations. Since $\theta_{k+1}$ is of dimension $\kappa$, which is greater than 3, it is necessary to use more than one location vector $r$ to solve the equation system. In general, $S=\lceil\kappa/3\rceil$ location vectors can be used to construct the equation system, i.e.,
$\underbrace{\begin{bmatrix}\Phi(r^{b_{k+1}}_{1})\\ \vdots\\ \Phi(r^{b_{k+1}}_{S})\\ \end{bmatrix}}_{\triangleq A}\theta_{k+1}=\underbrace{\begin{bmatrix}R_{b_{k}}^{b_{k+1}}\Phi(r^{b_{k}}_{1})\\ \vdots\\ R_{b_{k}}^{b_{k+1}}\Phi(r^{b_{k}}_{S})\\ \end{bmatrix}}_{\triangleq B(\psi_{k})}\theta_{k}.$ (10)
Note that the location vectors $r^{b_{k+1}}_{1},\cdots,r^{b_{k+1}}_{S}$ can be chosen such that $A$ has full column rank, and they are fixed. Finally, it holds that
@@
Since the magnetic field model should describe the magnetic field locally, the update in (11), which implicitly expands the applicable space $\Omega$ of the model, will inevitably introduce additional modeling errors. To account for those errors, with a slight abuse of notation, the update of the polynomial coefficients of the magnetic field as the body frame change is modeled as
$\theta_{k+1}=A^{\dagger}B(\psi_{k})\theta_{k}+w^{\theta}_{k},$ (12)
where $w_{k}^{\theta}$ is assumed to be a white Gaussian noise process with zero mean and covariance matrix $\Sigma_{\theta}$. Hence, the dynamic model for the magnetic-field subsystem is
@@
Note that $\psi_{k}$ is a function of $x^{\text{\tiny{ins}}}_{k}$, $\tilde{u}_{k}$, and $w^{\text{\tiny{ins}}}_{k}$.
II-C Magnetometer Array Measurement Model
@@ II-D Complete System
$H=\begin{bmatrix}0&\Phi(r_{m_{1}})\\ \vdots&\vdots\\ 0&\Phi(r_{m_{N}})\\ \end{bmatrix}.$ (17d)
@@ III State Estimation
The quaternion $q_{k}$ in the state vector does not belong to Euclidean space. Therefore, standard nonlinear filter algorithms, such as the extended Kalman filter and unscented Kalman filter, cannot be applied to the state-space model in (II-D) without appropriate modifications. The error state Kalman filter (ESKF) presented in [25] circumvents the problem by the use of “error quaternion”, which is a small perturbation around the “estimated quaternion” expressed in $\mathbb{R}^{3}$. With reference to Fig. 4, the algorithm works by propagating an estimated state $\hat{x}_{k}$ via the state transition model in (17a) and using a complementary Kalman filter to estimate the error state $\delta x_{k}$, which is then used to correct the estimated state $\hat{x}_{k}$. Since the ESKF is a standard algorithm, only the error propagation model that is unique to the state-space model in (II-D), will here be derived.
@@ III-C Magnetic Field Subsystem Error State Dynamics
$\begin{split}\Delta\hat{p}_{k}&=\hat{R}^{b_{k}}_{n}T_{s}(\hat{v}^{n}_{k}+g^{n}T_{s}/2)+\hat{s}_{k}T^{2}_{s}/2,\end{split}$ (23a)
@@
$\begin{split}\delta\Delta p_{k}=&\Delta p_{k}-\Delta\hat{p}_{k}\\ \approx&-[\epsilon_{k}]_{\times}\hat{R}^{b_{k}}_{n}T_{s}(\hat{v}^{n}_{k}+g^{n}T_{s}/2)+\hat{R}^{b_{k}}_{n}\delta v^{n}_{k}T_{s}\\ =&[\underbrace{\hat{R}^{b_{k}}_{n}T_{s}(\hat{v}^{n}_{k}+g^{n}T_{s}/2)}_{\eta(\hat{R}^{b_{k}}_{n},\hat{v}^{n}_{k})}]_{\times}\epsilon_{k}+\hat{R}^{b_{k}}_{n}\delta v^{n}_{k}T_{s}.\end{split}$ (24)
@@
$\delta\theta_{k+1}=F^{\theta}_{k}\delta\hat{x}_{k}+G^{\theta}_{k}w^{\theta}_{k},$ (26a)
@@
$G^{\theta}_{k}=\begin{bmatrix}0\quad-J_{2}T_{s}\quad I_{\kappa}\end{bmatrix}.$ (26d)
@@
Combining (III-B) and (III-C) gives the complete state-space description for the error state, i.e.,
@@
$F_{k}=\begin{bmatrix}F^{\text{\tiny{ins}}}_{k}&0\\ \hfil F^{\theta}_{k}\hfil\end{bmatrix},\quad G_{k}=\begin{bmatrix}G^{\text{\tiny{ins}}}_{k}\\ G^{\theta}_{k}\end{bmatrix}.$ (27c)

### 18 Mar 2024 v2

@@ II-B Magnetic Field Modeling
The magnetic field is a three-dimensional vector field whose properties are described by Maxwell’s equations. Let $M(r;\mu)$ be a model of the magnetic field at the location $r\in\mathbb{R}^{3}$, parameterized by the parameter $\mu$. When there is no free current in the space $\Omega$, $M(r;\mu)$ should fulfill
$\nabla_{r}\times M(r;\mu)$ $=0,$ (3a)
$\nabla_{r}\cdot M(r;\mu)$ $=0,$ (3b)
for all $r\in\Omega$[jackson1999Classical].
Equation (3a) guarantees that there exists a scalar potential function $\phi(r;\mu)$ whose gradient is $M(r;\mu)$, i.e.,
$M(r;\mu)=\nabla_{r}\phi(r;\mu).$ (4)
Let $\phi(r;\mu)$ be a polynomial of order $(l+1)$ in the $r$ component written as
$\phi(r;\mu)=h(r)^{\top}\mu+c.$ (5)
Here $h(r)$ is a vector whose elements are given by the product $r_{x}^{i}r_{y}^{j}r_{z}^{k}$ for $\forall i,j,k\in\mathbb{N}$, subject to $1\leq i+j+k\leq m$, $m=1,2,\cdots,l+1$. Further, $\mu\in\mathbb{R}^{L}$ is a column vector of dimension $L=(l+4)(l+3)(l+2)/6-1$ and $c$ is an arbitrary constant that does not affect the gradient of the potential function. Let $\Gamma(r)=\nabla_{r}h(r)^{\top}$, then $M(r;\mu)$ can be written as
$M(r;\mu)=\nabla_{r}\phi(r;\mu)=\Gamma(r)\mu.$ (6)
For the model $M(r;\mu)$ to fulfill (3b), the model parameters $\mu$ must be selected so that the following holds
$\nabla_{r}\cdot\Gamma(r)\mu=\sum_{i=x,y,z}\frac{d[\Gamma(r)\mu]_{i}}{dr_{i}}=0\quad\forall r\in\Omega.$ (7)
This constraint can be written as the linear equation system
$D\mu=0,$ (8)
where $D$ is a constant matrix derived in [23].
Equations (6) and (8) define the magnetic field model, which writes as
$M(r;\mu)=\Gamma(r)\mu,\quad D\mu=0.$ (9)
Finally, $M(r;\mu)$ can be reparameterized by introducing the matrix $D^{\perp}\triangleq\text{null}\{D\}$ whose columns span the null space of $D$, and then setting $\mu=D^{\perp}\theta$, where $\theta$ is a column vector of dimension equal to that of the null space of $D$. The reparameterized model is given by
$M(r;\theta)=\Phi(r)\theta,$ (10)
@@
where $\Phi(r)\triangleq\Gamma(r)D^{\perp}\in\mathbb{R}^{3\times\kappa}$ is the regression matrix defined in [23] and $\theta\in\mathbb{R}^{\kappa}$ is the coefficient of the polynomial model; for a $l^{\text{th}}$ order polynomial the model has $\kappa$ = dim($\theta$) = $l^{2}+4l+3$ unknown parameters[2]  [^1] An example of the 1st order model used in the experiment is provided in the appendix.. Note that the model (10) can be defined in either the body frame or navigation frame. Within this paper, it will be defined in the body frame. Next, a procedure for transforming the model from body frame $\alpha$ to body frame $\beta$ will be presented.
II-C Tranforming models between body frames
@@
Let the magnetic field model (10) be associated with the body frame, which means $M(r;\theta)$ accepts locations $r$ expressed in the current body frame and outputs magnetic field vector in the same frame. Then the magnetic field can be represented in the two body frames $\alpha$ and $\beta$, i.e.,
$\begin{split}M^{\alpha}(r^{\alpha};\theta^{\alpha})&=\Phi(r^{\alpha})\theta^{\alpha},\\ M^{\beta}(r^{\beta};\theta^{\beta})&=\Phi(r^{\beta})\theta^{\beta}.\\ \end{split}$ (11)
Here the superscripts on $M$ and $r$ denote the corresponding body frame in which they are resolved and the superscript on $\theta$ denotes the frame with which the model coefficients are associated.
By expressing the magnetic field vector at a given location with the two models and aligning them in the same frame, the two models can be related as
$M^{\beta}(r^{\beta};\theta^{\beta})=R_{\alpha}^{\beta}M^{\alpha}(r^{\alpha};\theta^{\alpha}),$ (12a)
where
$r^{\alpha}=(R_{\alpha}^{\beta})^{\top}r^{\beta}+\Delta p^{\alpha}.$ (12b)
Here $\Delta p^{\alpha}$ denotes the translation vector expressed in body frame $\alpha$. An illustration of the geometric relationship between the two frames and the magnetic field vector is shown in Fig. 3.
Now consider transforming the magnetic field model from body frame $b_{k}$ to $b_{k+1}$. Let the relative body frame change be encoded by
$\psi_{k}=\begin{bmatrix}\Delta p_{k}^{b_{k}}\\ \Delta\phi_{k}\end{bmatrix},$ (13)
where ${\Delta p}_{k}^{b_{k}}\in\mathbb{R}^{3}$ and $\Delta\phi_{k}\in[0,2\pi]^{3}$ denote the translation and orientation change from the body frame $b_{k}$ to $b_{k+1}$, respectively. Replacing $\alpha$ and $\beta$ with $b_{k}$ and $b_{k+1}$ respectively and using $\theta_{k}$ and $\theta_{k+1}$ to denote the corresponding coefficients, the following holds
$M^{b_{k+1}}(r^{b_{k+1}};\theta_{k+1})=R_{b_{k}}^{b_{k+1}}M^{b_{k}}(r^{b_{k}};\theta_{k}),$ (14a)
where
$r^{b_{k}}=(R_{b_{k}}^{b_{k+1}})^{\top}r^{b_{k+1}}+{\Delta p}_{k}^{b_{k}}.$ (14b)
The rotation matrix $R_{b_{k}}^{b_{k+1}}$ and translation $\Delta p_{k}^{b_{k}}$ are given by
$R_{b_{k}}^{b_{k+1}}=\left(\text{exp}(\left[\Delta\phi_{k}\right]_{\times})\right)^{\top},$ (15a)
$\Delta p_{k}^{b_{k}}=R_{b_{k}}^{n^{\top}}\left(v_{k}^{n}T_{s}+(R^{n}_{b_{k}}s_{k}+g^{n})\frac{T_{s}^{2}}{2}\right).$ (15b)
@@
Here $\Delta\phi_{k}=\omega_{k}T_{s}$ and $[\cdot]_{\times}$ is an operator that maps a vector in $\mathbb{R}^{3}$ to a skew-symmetric matrix such that $\left[a\right]_{\times}b=a\times b$.
[…]
@@
Figure 3: A 2D illustration of the geometric relationship between the body frames at two consecutive times. The applicable region $\Omega$ of the magnetic field model at time $k$ is in blue, and the black dot indicates the location where the two models output the corresponding magnetic field in their coordinate frames.
@@
Next, substituting the generic magnetic model in (II-C) with the proposed polynomial model (10) yields the equality
@@
Note that for a given $\{r^{b_{k+1}},\theta_{k},\psi_{k}\}$, (16) represents 3 linear equations. Since $\theta_{k+1}$ is of dimension $\kappa$, which is greater than 3, it is necessary to use more than one location vector $r$ to solve the equation system. In general, $S=\lceil\kappa/3\rceil$ location vectors can be used to construct the equation system
$A\theta_{k+1}=B(\psi_{k})\theta_{k},$ (17a)
where
$A=\begin{bmatrix}\Phi(r^{b_{k+1}}_{1})\\ \vdots\\ \Phi(r^{b_{k+1}}_{S})\\ \end{bmatrix},\quad B(\psi_{k})=\begin{bmatrix}R_{b_{k}}^{b_{k+1}}\Phi(r^{b_{k}}_{1})\\ \vdots\\ R_{b_{k}}^{b_{k+1}}\Phi(r^{b_{k}}_{S})\\ \end{bmatrix}.$ (17b)
If the vectors $r^{b_{k+1}}_{1},\cdots,r^{b_{k+1}}_{S}$ can be chosen such that $A$ has full column rank, it holds that
@@
Since the magnetic field model should describe the magnetic field locally, the update in (18), which with time implicitly expands the applicable space $\Omega$ of the model, will inevitably introduce modeling errors. To account for those errors, with a slight abuse of notation, the update of the polynomial coefficients of the magnetic field as the body frame change is modeled as
@@
Here $w_{k}^{\theta}$ is assumed to be a white Gaussian noise process with zero mean and covariance matrix $\Sigma_{\theta}$. Note that $\psi_{k}$ is a function of $x^{\text{\tiny{ins}}}_{k}$, $\tilde{u}_{k}$, and $w^{\text{\tiny{ins}}}_{k}$.
II-D Magnetometer Array Measurement Model
@@ II-E Complete System
$H=\begin{bmatrix}0_{3\times 16}&\Phi(r_{m_{1}})\\ \vdots&\vdots\\ 0_{3\times 16}&\Phi(r_{m_{N}})\\ \end{bmatrix}.$ (23d)
@@ III State Estimation
The quaternion $q_{k}$ in the state vector does not belong to Euclidean space. Therefore, standard nonlinear filter algorithms, such as the extended Kalman filter and unscented Kalman filter, cannot be applied to the state-space model in (II-E) without appropriate modifications. The error state Kalman filter (ESKF) presented in [25] circumvents the problem by the use of “error quaternion”, which is a small perturbation around the “estimated quaternion” expressed in $\mathbb{R}^{3}$. With reference to Fig. 4, the algorithm works by propagating an estimated state $\hat{x}_{k}$ via the state transition model in (23a) and using a complementary Kalman filter to estimate the error state $\delta x_{k}$, which is then used to correct the estimated state $\hat{x}_{k}$. Since the ESKF is a standard algorithm, only the error propagation model that is unique to the state-space model in (II-E), will here be derived.
@@ III-C Magnetic Field Subsystem Error State Dynamics
$\begin{split}\Delta\hat{p}_{k}^{b_{k}}&=\hat{R}^{b_{k}}_{n}T_{s}(\hat{v}^{n}_{k}+g^{n}T_{s}/2)+\hat{s}_{k}T^{2}_{s}/2,\end{split}$ (29a)
@@
$\begin{split}\delta\Delta p_{k}=&\Delta p_{k}^{b_{k}}-\Delta\hat{p}_{k}^{b_{k}}\\ \approx&-[\epsilon_{k}]_{\times}\hat{R}^{b_{k}}_{n}T_{s}(\hat{v}^{n}_{k}+g^{n}T_{s}/2)+\hat{R}^{b_{k}}_{n}\delta v^{n}_{k}T_{s}\\ =&[\underbrace{\hat{R}^{b_{k}}_{n}T_{s}(\hat{v}^{n}_{k}+g^{n}T_{s}/2)}_{\eta(\hat{R}^{b_{k}}_{n},\hat{v}^{n}_{k})}]_{\times}\epsilon_{k}+\hat{R}^{b_{k}}_{n}\delta v^{n}_{k}T_{s}.\end{split}$ (30)
@@
$\delta\theta_{k+1}=F^{\theta}_{k}\delta\hat{x}_{k}+G^{\theta}_{k}\begin{bmatrix}w^{\text{\tiny{ins}}}_{k}\\ w^{\theta}_{k}\end{bmatrix},$ (32a)
@@
$G^{\theta}_{k}=\begin{bmatrix}0\quad-A^{\dagger}J_{2}T_{s}\quad 0\quad 0\quad I_{\kappa}\end{bmatrix}.$ (32d)
@@
Combining (III-B) and (III-C) gives the complete state-space model of the error state, i.e.,
@@
$F_{k}=\begin{bmatrix}F^{\text{\tiny{ins}}}_{k}&0\\ \hfil F^{\theta}_{k}\hfil\end{bmatrix},\quad G_{k}=\begin{bmatrix}G^{\text{\tiny{ins}}}_{k}&0\\ \hfil G^{\theta}_{k}\hfil\end{bmatrix},$ (33c)
$H_{\delta x}=\begin{bmatrix}0_{3\times 15}&\Phi(r_{m_{1}})\\ \vdots&\vdots\\ 0_{3\times 15}&\Phi(r_{m_{N}})\\ \end{bmatrix}.$ (33g)
Here $\delta y_{k}\triangleq y_{k}-H\hat{x}_{k}$.